### v2

# The minimal periodicity for integral bases of pure number fields

Abstract.
Fix $n\geq 3$. For the pure field $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$, where $a\neq\pm 1$ is $n$th-power-free, we encode an integral basis in the fixed coordinate $\{1,\theta,\dots,\theta^{n-1}\}$ by its shape. We prove a sharp local-to-global principle: for each $p^{e}\!\parallel n$, the local shape at $p$ is determined by $a\bmod p^{\,e+1}$, and this precision is optimal. Moreover, the global shape is periodic with minimal modulus
$M(n)=\prod_{p^{e}\parallel n}p^{\,e+1}=n\cdot\mathrm{rad}(n),$
providing many applications in the understanding integral bases of pure number fields.
MSC 2020. 11R04, 11R21.
Keywords. Integral basis, pure fields, periodicity, monogeneity.
1. Introduction
A classical problem in algebraic number theory is to study the ring of integers $\mathcal{O}_{K}$ of a number field $K$, in particular for the pure fields $K=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ (see [GAÁ19] for a modern treatment). Throughout we assume that $a$ is $n$th-power-free, $a\neq\pm 1$, and that $x^{n}-a$ is irreducible over $\mathbb{Q}$ (equivalently, $K_{a}:\mathbb{Q}]=n$). Recently, two complementary phenomena have guided works on this direction. First, in small degrees ($3\leq n\leq 9$) Gaál–Remete [GR17] showed that the global pattern of integral bases is periodic as $a$ varies, with period $n^{2}$. In particular, if one writes an integral basis with rational coefficients in the $\{1,\theta,\dots,\theta^{n-1}\}$-coordinates, then every prime dividing a denominator must divide $n$ (a global restriction that already foreshadows a local explanation). This offers a striking, hands‑on picture of regularity in families of pure fields. Secondly, for general $n$ and squarefree $a$, Jakhar–Khanduja–Sangwan [JKS21] gave explicit integral bases and proved a global period $n\cdot\prod_{p\mid n}p$. Their construction is intrinsically local: it splices together $p$-integral bases at primes $p\mid n$ and encodes, for each $m=1,\dots,n-1$, a pp‑adic denominator exponent and a small “linear correction”. The dependence on $a$ is controlled by a short list of pp‑adic invariants (e.g. $v_{p}(a^{p-1}-1)$), revealing that only one extra pp‑adic digit beyond $p^{e_{p}}\parallel n$ is ever needed to fix the local shape at $p$.
In this paper we answer the question on how much information about $a$ do the local computations at each $p\mid n$ really need? Equivalently, what is the smallest modulus $M(n)$ such that, whenever $a\equiv a^{\prime}\pmod{M(n)}$, the structure of the integral basis of $\mathcal{O}_{\mathbb{Q}(\sqrt[n]{a})}$ is the same as for $a^{\prime}$? We show that the Jakhar–Khanduja–Sangwan period is in fact minimal, in a precise sense, for all $n\geq 3$, and we do so in a way that emphasizes the underlying local mechanism. The first novelty of our paper is we unify and formalize these two phenomena by introducing the concept of shape (defined precisely in  2.2) as the sequence, for $m=1,\dots,n-1$, consisting of (i) the prime‐power denominators supported on $p\mid n$ and (ii) the corresponding linear corrections recorded modulo those denominators. We then prove that (see Theorem 3.1) the smallest global modulus governing periodicity of the shape is always
$M(n)=\prod_{p^{e_{p}}\parallel n}p^{\,e_{p}+1}\;=\;n\cdot\mathrm{rad}(n)$
In particular $M(n)\mid n^{2}$ for all $n$, with equality $M(n)=n^{2}$ if and only if $n$ is squarefree. Thus we both recover the $n^{2}$-period of [GR17] and pinpoint exactly when it is sharp.
Method
Our proof is short and local-to-global and uses only the explicit $p$-local invariants and bases of Jakhar–Khanduja–Sangwan [JKS21] under Hypothesis (H). Fix a prime power $p^{e}\parallel n$ and write $v_{p}$ for the $p$-adic valuation. When $v_{p}(a)=0$ we set
$r_{p}(a)=v_{p}(a^{p-1}-1)-1,\qquad d_{p}(a)=\min\{r_{p}(a),e\},$
and when $v_{p}(a)>0$ we have $r_{p}(a)=-1$ (hence $d_{p}(a)=-1$). In [JKS21, Theorem 1.6], the pp–denominator exponents $k_{p,m}$ are explicit functions of $p,n,m$ and the truncated invariant $d_{p}(a)=\min\{v_{p}(a^{p-1}-1)-1,e\}$ (in the unit branch $v_{p}(a)=0$). Moreover, the correction terms $\beta_{m}$ are constructed from congruence data modulo $p^{\,k_{p,m}+1}$ (in particular via an inverse of $a$ modulo $p^{\,k_{p,m}+1}$; see the construction preceding [JKS21, Theorem 1.6]). Hence the reduced classes $\beta_{m}\bmod p^{\,k_{p,m}}$ are determined by $a\bmod p^{e+1}$, since $k_{p,m}\leq e$. Since $d_{p}(a)$ is determined by $a\bmod p^{e+1}$, the local $p$-shape is determined by $a\bmod p^{e+1}$.
Already among $p$-adic units one can choose
$a\equiv 1+p^{e}u\not\equiv 1\pmod{p^{e+1}}\qquad\text{and}\qquad a^{\prime}\equiv 1+p^{e+1}u^{\prime}\pmod{p^{e+2}}$,
with $p\nmid uu^{\prime}$, for which $r_{p}(a)=e-1$ while $r_{p}(a^{\prime})\geq e$, hence $d_{p}(a)\neq d_{p}(a^{\prime})$. Therefore the $p$-part of the shape changes, so $p^{e+1}$ is the least local modulus at $p$.
By the Chinese Remainder Theorem the global shape depends only on $a\bmod\prod_{p\mid n}p^{e_{p}+1}=n\cdot\mathrm{rad}(n)$, giving the optimal upper bound; local sharpness at every $p\mid n$ makes this modulus minimal.
Applications and further consequences
The point of view developed here—the local shape at each $p\mid n$ and its global gluing—yields several immediate applications. Firstly, we can show that one can compute, for any fixed $n$, a finite lookup table indexed by the classes $a\bmod M(n)$ recording the pp–denominator exponents and reduced corrections for every $p\mid n$. Thanks to Theorem 3.1, the table itself can be filled entirely locally using the explicit local recipes in [JKS21]. This offers a uniform and predictable alternative to ad hoc casework used in small degrees.
The second application concerns local tests for the $p$-part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ at primes $p\mid n$. When $p\parallel n$ and $v_{p}(a)=0$, the dichotomy between $a^{p-1}\not\equiv 1\pmod{p^{2}}$ and $a^{p-1}\equiv 1\pmod{p^{2}}$ decides whether $p$ divides $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ (equivalently, whether $\mathbb{Z}[\theta]$ is pp–maximal). More generally, for $p^{e}\parallel n$ and $v_{p}(a)=0$ the relevant invariant is $d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,e\,\}$, and Theorem 3.1 shows that this (and with it all pp–denominator exponents and reduced corrections in an explicit basis) is determined by $a\bmod p^{e+1}$ and cannot be recovered from $a\bmod p^{e}$ alone.
Finally, because the shape is a class function of $a\bmod M(n)$, families of fields cut out by finitely many local conditions on $a\bmod M(n)$ admit explicit counting and density statements. Using the equidistribution of $n$th–power–free integers in arithmetic progressions, one obtains asymptotics and natural densities for any prescribed union of shape classes; in particular, the global density factorises into a product of local proportions attached to the sets of admissible residues in $(\mathbb{Z}/p^{e_{p}+1}\mathbb{Z})^{\times}$.
Historical remarks and related results
The study of integral bases and monogeneity goes back to Dedekind and Hensel and is tightly intertwined with the index, discriminant, and the factorization of rational primes in $\mathcal{O}_{K}$. Dedekind’s criterion and index theorem explain how, when $p\nmid[\mathcal{O}_{K}:\mathbb{Z}[\theta]]$, the prime-ideal factorization of $p\mathcal{O}_{K}$ is read off from the factorization of the minimal polynomial modulo $p$ (see, e.g., [DED78]). Hensel developed local methods that relate common index divisors to the splitting pattern of $p$ [HEN08].
When Dedekind’s criterion fails (i.e. some $p\mid[\mathcal{O}_{K}:\mathbb{Z}[\theta]]$), the modern local machine is based on Newton polygons at $p$, initiated by Ore in the 1920s and refined into a systematic algorithm by Montes and collaborators: one builds (higher-order) $p$-adic Newton polygons, reads residual polynomials along principal sides, and from this recovers $p$-adic factorization, indices, discriminants, and $p$-integral bases [ORE28, MN92, GMN12, GMN15]. This underlies much current algorithmic number theory around integral bases and prime decomposition. For practical $p$-integral basis extraction and index computations using Newton polygons, see ElFadil and his collaborators’ papers [EMN12, EL 20] and expositions cited therein. For the complexity and implementation discussions, take a look at [BNS13]. These provide worked implementations that agree with and conceptually support the $p$-local pieces used by Gaál–Remete and Jakhar–Khanduja–Sangwan.
In Kummer or radical extensions over cyclotomic bases, the decisive local congruences at tame primes naturally live one level deeper in the local parameter: modulo $(1-\zeta_{p})^{2}$ (the uniformizer at $p$ in $\mathbb{Z}[\zeta_{p}]$), mirroring the pure-field “one more $p$-adic digit” from $p$ to $p^{2}$. For a modern treatment proving relative monogeneity criteria and non-monogeneity results in this framework see Smith [SMI19, SMI21].
Organization of the paper
Section 2 introduces the shape of an integral basis and gives illustrative examples. Section 3 is preceded by our main local–to–global result (Theorem 3.1) and its direct consequences. We also include a tame relative version (Proposition 3.6) and examples. Section 4 develops applications: sharp local power-basis, explicit discriminant jumps, and counting/density results for prescribed shape classes, with worked quartic and sextic cases.
Acknowledgments
We are grateful to the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment. We really appreciate the anonymous reviewer for their careful reading of the manuscript and for their valuable comments and suggestions. We thank István Gaál for his interest in our work and for his helpful comments.
Conflict of Interest
The author declares that there is no conflict of interest.
Data availability
This research is purely theoretical; no datasets were generated or analyzed. All mathematical results and worked examples are contained in the article.
2. Shape, periodicity, and minimal period
Throughout let $n\geq 3$ and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$, where $a\in\mathbb{Z}$ is $n$th‑power‑free, $a\neq\pm 1$, and $x^{n}-a$ is irreducible over $\mathbb{Q}$. Write $n=\prod_{p}p^{e_{p}}$ and $\mathrm{rad}(n)=\prod_{p\mid n}p$. When we write an integral basis in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates we mean a $\mathbb{Z}$–basis of the form
{ $\Bigl\{\,1,\ \frac{\theta^{m}+\beta_{m}}{D_{m}}\ :\ 1\leq m\leq n-1\Bigr\},$
Here $D_{m}$ denotes the (normalized) denominator of the $m$th basis element in these fixed coordinates, after clearing the factors supported on primes dividing $a$ into the numerator (cf. Remark 2.4), with $D_{m}\in\mathbb{Z}_{\geq 1}$, $\beta_{m}\in\mathbb{Z}[\theta]$ and linear in $\{1,\theta,\dots,\theta^{m-1}\}$. We use the following standard assumption in [JKS21] under which the explicit integral bases of $\mathcal{O}_{K_{a}}$ are stated. In particular, it holds whenever $a$ is squarefree or $(a,n)=1$.
Hypothesis (H): We say that the pair $(n,a)$ satisfies (H) if $a$ is $n$th–power–free and, for every prime $p\mid n$, either $v_{p}(a)=0$ or the exponent $v_{p}(a)$ is not divisible by $p$ (i.e. $p\nmid v_{p}(a)$).
Write $n=\prod_{p}p^{e_{p}}$ and factor $a$ (uniquely up to sign) as
$a=\pm\prod_{j=1}^{n-1}a_{j}^{\,j},$
with the $a_{j}$ squarefree and pairwise coprime. For $0\leq m\leq n-1$ set
$C_{m}(a)\ :=\ \prod_{j=1}^{n-1}a_{j}^{\lfloor jm/n\rfloor}.$
For each $p\mid n$ define
$r_{p}(a)\ :=\ \begin{cases}v_{p}(a^{p-1}-1)-1,&v_{p}(a)=0,\\[2.0pt] -1,&v_{p}(a)>0,\end{cases}\qquad d_{p}(a)\ :=\ \min\{\,r_{p}(a),\,e_{p}\,\}.$
For each $m$ and each $p\mid n$, let $k_{p,m}$ be the largest integer $k$ with $0\leq k\leq d_{p}(a)$ such that $m\geq n-n/p^{\,k}$ (put $k_{p,m}=0$ if no such $k$ exists). We obtain the following remarkable result, which is a reformulation of [JKS21, Theorem 1.6]. The correction terms $\beta_{m}$ (and hence their reductions modulo $p^{k_{p,m}}$) are given explicitly by the construction preceding [JKS21, Theorem 1.6], based on the auxiliary elements $\delta_{i,m}$ from [JKS21, Lemma 1.2]; see also [JKS21, Remark 1.7] for the freedom in choosing inverses.
Theorem 2.1 (Explicit local/global form of an integral basis).
Assume (H). Then there exist elements $\beta_{m}\in\mathbb{Z}[\theta]$, linear in $\{1,\theta,\dots,\theta^{m-1}\}$, such that
:={ $\mathcal{B}\ :=\ \Bigl\{\,1,\ \frac{\theta^{m}+\beta_{m}}{\;C_{m}(a)\cdot\prod_{p\mid n}p^{\,k_{p,m}}\;}\ :\ 1\leq m\leq n-1\Bigr\}$
is an integral basis of $\mathcal{O}_{K_{a}}$. Moreover, for each $p\mid n$ the residue class $\beta_{m}\bmod p^{\,k_{p,m}}$ is well determined by $p,n,m$ and $a$ via the construction, and $C_{m}(a)$ is supported only at primes dividing $a$.
Fix $p\mid n$. For each $1\leq m\leq n-1$ let $k_{p,m}\geq 0$ be the exponent of $p$ in the denominator of the $m$th basis element when an integral basis is written in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates. In the explicit bases of Theorem 2.1 a linear correction $\beta_{m}$ appears; we record its reduction modulo the full pp–power denominator. We formalize the insight in the following notion.
Definition 2.2 (Local and global shape).
For $p\mid n$, the local pp–shape of $\mathcal{O}_{K_{a}}$ is given by
$\mathbf{S}_{p}(a)\ :=\ \Bigl((k_{p,m})_{m=1}^{n-1},\ (\beta_{m}\!\!\!\!\pmod{p^{k_{p,m}}})_{m=1}^{n-1}\Bigr).$
The shape of $\mathcal{O}_{K_{a}}$ is the tuple of local shapes
$\mathbf{S}(a):=(\mathbf{S}_{p}(a))_{p\mid n}.$
Remark 2.3 (Convention for $k_{p,m}=0$).
If $k_{p,m}=0$, then we regard the residue class $\beta_{m}\bmod p^{\,k_{p,m}}$ as trivial; in particular the pp–local shape depends only on those $\beta_{m}$ with $k_{p,m}>0$.
Following [JKS21], we ignore the factors $C_{m}(a)$ coming solely from primes dividing $a$.
Remark 2.4 (No denominator primes away from $n$ in the normalized form).
Write an integral basis in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates in the normalized manner of Definition 2.2, i.e. with all factors supported on primes dividing $a$ absorbed into the numerators $C_{m}(a)$. Then every prime divisor of each denominator $D_{m}$ divides $n$. Indeed, in the explicit bases of Jakhar–Khanduja–Sangwan the denominators are exactly $\prod_{p\mid n}p^{k_{p,m}}$, while all primes dividing $a$ appear only in $C_{m}(a)$. Hence, after normalizing by clearing the $C_{m}(a)$ into the numerators, no prime $q\nmid n$ can appear in any $D_{m}$.
Definition 2.5 (Periodicity and minimal period).
Fix $n$ and let $N\geq 1$. We say that the shape is periodic modulo $N$ if, for all $n$th–power–free integers $a,a^{\prime}\in\mathbb{Z}$ for which $\mathbf{S}(a)$ and $\mathbf{S}(a^{\prime})$ are defined,
$a\equiv a^{\prime}\pmod{N}\qquad\Longrightarrow\qquad\mathbf{S}(a)=\mathbf{S}(a^{\prime}).$
A positive integer $N$ with this property is called a period (for fixed $n$). The minimal period is the least such $N$.
We will show that the shape is periodic modulo
[…]
and the local sharpness will force $M(n)$ is minimal.
Example 2.6 (Small degrees $3\leq n\leq 9$).
Formula (1) gives
M(3)=9,M(4)=8,M(5)=25,M(6)=36,M(7)=49,M(8)=16,M(9)=27.M(3)=9,\ M(4)=8,\ M(5)=25,\ M(6)=36,\ M(7)=49,\ M(8)=16,\ M(9)=27.
Thus Gaál–Remete’s $n^{2}$–periodicity for $3\leq n\leq 9$ is minimal iff $n$ is squarefree; otherwise $M(n)<n^{2}$, in agreement with their observation.
We record two examples that instantiate the definitions of the local and global shape and line up with known tables.
Example 2.7 (Quartic case $n=4$).
Assume Hypothesis (H). Let $a$ be 4th‑power‑free. Put $r=v_{2}(a-1)-1$ when $a$ is odd, and r=1r=-1 when $2\mid a$. Then $d=\min\{r,2\}$. The general rule above gives
$k_{2,1}=0,\qquad k_{2,2}=\mathbf{1}_{\{d\geq 1\}},\qquad k_{2,3}=\begin{cases}2,&d=2,\\ 1,&d=1,\\ 0,&d=0.\end{cases}$
So there are exactly three 2‑adic denominator patterns among odd $a$, according to $a\bmod 8$ (and no 2‑denominators when $2\mid a$). Using Gaál–Remete’s explicit bases, one can choose the accompanying corrections $\beta_{m}$ (reduced modulo the full 2‑power denominator) as follows :
Quartic case $n=4$: basis and corrections (mod $8$).
$a\bmod 8$ $(k_{2,1},k_{2,2},k_{2,3})$ basis fragment reduced corrections
$3,7$ $(0,0,0)$ { $\{\,1,\theta,\theta^{2},\theta^{3}\,\}$ $\beta_{2}\equiv 0,\ \ \beta_{3}\equiv 0\ (\bmod 1)$
55 $(0,1,1)$ { $\bigl\{\,1,\theta,\frac{\theta^{2}+1}{2},\frac{\theta^{3}+\theta}{2}\,\bigr\}$ $\beta_{2}\equiv 1\ (\bmod 2),\ \ \beta_{3}\equiv\theta\ (\bmod 2)$
11 $(0,1,2)$ { $\bigl\{\,1,\theta,\frac{\theta^{2}+1}{2},\frac{\theta^{3}+\theta+\theta^{2}+1}{4}\,\bigr\}$ $\beta_{2}\equiv 1\ (\bmod 2),\ \ \beta_{3}\equiv 1+\theta+\theta^{2}\ (\bmod 4)$
$2,6$ $(0,0,0)$ { $\{\,1,\theta,\theta^{2},\theta^{3}\,\}$ $\beta_{2}\equiv 0,\ \ \beta_{3}\equiv 0\ (\bmod 1)$
Local shape $\mathbf{S}_{2}(a)$ is completely determined by $a\bmod 8$ (this is the minimal local modulus $2^{\,e_{2}+1}$).Global shape equals the local one here (since $2$ is the only denominator prime). These three patterns recur with period $2^{e_{2}+1}=8$, in agreement with both our theorem and the quartic tables in [GR17].
Example 2.8 (Sextic case $n=6$).
Assume Hypothesis (H) (in particular $(a,6)=1$ in the tables below). Here $p=2,3$ with $e_{2}=e_{3}=1$ and $M(6)=2^{2}3^{2}=36$. The local moduli are $4$ at $p=2$ and $9$ at $p=3$. We tabulate the local shapes and then the global CRT gluing.
Local table at $p=2$ (mod $4$).
$a$ (mod $4$) $(k_{2,1},k_{2,2},k_{2,3},k_{2,4},k_{2,5})$ 2-part denoms reduced $2$-adic $\beta$’s
11 $(0,0,1,1,1)$ m=3,4,5: 2m=3,4,5:\ 2 $\beta_{3}\equiv 1,\ \ \beta_{4}\equiv\theta,\ \ \beta_{5}\equiv\theta^{2}\ (\mathrm{mod}\ 2)$
$0,2,3$ $(0,0,0,0,0)$ none trivial
Local table at $p=3$ (mod $9$).
$a$ (mod $9$) $(k_{3,1},k_{3,2},k_{3,3},k_{3,4},k_{3,5})$ 3-part denoms reduced $3$-adic $\beta$’s
11 $(0,0,0,1,1)$ m=4,5: 3m=4,5:\ 3 $\beta_{4}\equiv 1+\theta^{2},\ \ \beta_{5}\equiv\theta+\theta^{3}\ (\mathrm{mod}\ 3)$
88 $(0,0,0,1,1)$ m=4,5: 3m=4,5:\ 3 $\beta_{4}\equiv 1+2\theta^{2},\ \ \beta_{5}\equiv\theta+2\theta^{3}\ (\mathrm{mod}\ 3)$
$0,2,3,4,5,6,7$ $(0,0,0,0,0)$ none trivial
Global (CRT) gluing modulo $36$.
Congruence on $a$ $\mathbf{(k_{2,m})}$ $\mathbf{(k_{3,m})}$ Consequence
$a\equiv 1\ (\mathrm{mod}\ 4),\ a\equiv\pm 1\ (\mathrm{mod}\ 9)$ $k_{2,3}=k_{2,4}=k_{2,5}=1$ $k_{3,4}=k_{3,5}=1$ $m=4,5$ have denom. $2\cdot 3=6$
$a\equiv 1\ (\mathrm{mod}\ 4),\ a\not\equiv\pm 1\ (\mathrm{mod}\ 9)$ $k_{2,3}=k_{2,4}=k_{2,5}=1$ all 0 only $2$-adic denominators at $m=3,4,5$
$a\not\equiv 1\ (\mathrm{mod}\ 4),\ a\equiv\pm 1\ (\mathrm{mod}\ 9)$ all 0 $k_{3,4}=k_{3,5}=1$ only $3$-adic denominators at $m=4,5$
otherwise all 0 all 0 no denominators
All data are determined by a(mod 36)a\ ($\mathrm${mod}\ 36); the $\beta$‑reductions are taken modulo the full pp‑power denominator in each row.
3. Main theorems
In this section we prove the precise local-to-global statement that underlies the paper: at each prime $p^{e}\parallel n$ the local pp–shape of $\mathcal{O}_{K_{a}}$ is decided by a single additional pp‑adic digit beyond $p^{e}$, and this is best possible; consequently the global shape is periodic with minimal modulus
$M(n)\;=\;\prod_{p^{e}\parallel n}p^{\,e+1}\;=\;n\cdot\operatorname{rad}(n).$
As a preliminary reduction, if an integral basis is written in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates, by the Gaál–Remete theorem, only primes dividing $n$ can occur in the denominators in the chosen coordinates, so the local shape $\mathbf{S}_{p}(a)$ is only defined and only varies for $p\mid n$.
The proof of our main theorem uses only the explicit local invariants and bases in Jakhar–Khanduja–Sangwan [JKS21] (under Hypothesis (H)) and a Chinese remainder argument.
We now prove our main result.
Theorem 3.1 (Minimal local/global modulus for the shape).
Let $n\geq 3$, $a\in\mathbb{Z}$ be $n$th–power–free with $a\neq\pm 1$, and assume $x^{n}-a$ is irreducible over $\mathbb{Q}$, and $K_{a}=\mathbb{Q}(\theta)$ with $\theta^{n}=a$ satisfying hypothesis (H). For each prime $p\mid n$ write $p^{e_{p}}\parallel n$ and let $\mathbf{S}_{p}(a)$ and $\mathbf{S}(a)$ be the local and global shapes.
  1. (i)
    (Local determinacy.) For every $p\mid n$, the local shape $\mathbf{S}_{p}(a)$ depends only on the residue class of $a$ modulo $p^{\,e_{p}+1}$. Equivalently, if $a^{\prime}$ is another $n$th–power–free integer satisfying the same standing hypotheses (in particular (H) and irreducibility) and $a\equiv a^{\prime}\pmod{p^{\,e_{p}+1}}$, then $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
  2. (ii)
    (Local sharpness.) For every $p\mid n$ there exist $n$th–power–free integers $a,a^{\prime}$ with $a\equiv a^{\prime}\pmod{p^{e_{p}}}$ but $a\not\equiv a^{\prime}\pmod{p^{e_{p}+1}}$ such that $\mathbf{S}_{p}(a)\neq\mathbf{S}_{p}(a^{\prime})$. Thus $p^{e_{p}}$ precision does not suffice in general.
  3. (iii)
    (Global minimal period.) The global shape $\mathbf{S}(a)$ is periodic modulo
    $M(n)\;:=\;\prod_{p^{e_{p}}\parallel n}p^{\,e_{p}+1}=n\cdot\operatorname{rad}(n),$
    and $M(n)$ is the least such modulus: if $N\geq 1$ has the property that for all $n$th–power–free $a,a^{\prime}$ satisfying the standing hypotheses,
    $a\equiv a^{\prime}\pmod{N}\qquad\Longrightarrow\qquad\mathbf{S}(a)=\mathbf{S}(a^{\prime}),$
    we get $M(n)\mid N$.
Proof.
Fix a prime $p\mid n$ with $p^{e}\parallel n$ and write $v=v_{p}$. Recall the invariants
$r_{p}(a)=\begin{cases}v_{p}(a^{p-1}-1)-1,&v_{p}(a)=0,\\ -1,&v_{p}(a)>0,\end{cases}\qquad d_{p}(a)=\min\{r_{p}(a),e\},$
and the denominator exponents $k_{p,m}$ defined in §2. By [JKS21, Theorem 1.6 and Theorem 1.8] (under hypothesis (H)), the pp–denominator exponents $k_{p,m}$ and the reduced correction terms $\beta_{m}\bmod p^{k_{p,m}}$ in the basis of Theorem 2.1 are determined by $p,n,m$ and the single invariant $d_{p}(a)$ (together with the branch $v_{p}(a)=0$ or $v_{p}(a)>0$).
(i) Local determinacy. Assume $a\equiv a^{\prime}\pmod{p^{e+1}}$. Then $v_{p}(a)>0$ if and only if $v_{p}(a^{\prime})>0$. If $v_{p}(a)>0$, then $r_{p}(a)=r_{p}(a^{\prime})=-1$ and hence $d_{p}(a)=d_{p}(a^{\prime})=-1$; therefore $k_{p,m}=0$ for all $m$. By Remark 2.3 the reduced correction data are trivial, so $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
Assume now $v_{p}(a)=0$ (hence also $v_{p}(a^{\prime})=0$). From $a\equiv a^{\prime}\pmod{p^{e+1}}$ we have $a^{p-1}\equiv(a^{\prime})^{p-1}\pmod{p^{e+1}}$. Therefore
$\min\bigl\{v_{p}(a^{p-1}-1),\,e+1\bigr\}=\min\bigl\{v_{p}((a^{\prime})^{p-1}-1),\,e+1\bigr\}.$
It follows that
$d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,e\,\}=\min\{\,v_{p}((a^{\prime})^{p-1}-1)-1,\,e\,\}=d_{p}(a^{\prime}).$
Since the exponents $k_{p,m}$ are defined purely from $d_{p}(a)$ (together with $p,n,m$), we get $(k_{p,m})_{m=1}^{n-1}=(k^{\prime}_{p,m})_{m=1}^{n-1}$. Finally, in the construction underlying [JKS21, Theorem 1.6], each reduced class $\beta_{m}\bmod p^{\,k_{p,m}}$ is computed from congruence data modulo $p^{\,k_{p,m}+1}$ (in particular from an inverse of $a$ modulo $p^{\,k_{p,m}+1}$). Since $k_{p,m}\leq e$ and $a\equiv a^{\prime}\pmod{p^{e+1}}$, the required congruence data agree for $a$ and $a^{\prime}$, hence the reduced correction terms coincide. Therefore $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$.
(ii) Local sharpness. Take $p$-adic units
$a=1+p^{e}u,\qquad a^{\prime}=1+p^{e+1}u^{\prime},\qquad p\nmid uu^{\prime}.$
By replacing $a$ and $a^{\prime}$ by congruent integers if necessary, we may (and do) assume that $a$ and $a^{\prime}$ are squarefree and coprime to $n$; in particular they are $n$th–power–free, satisfy hypothesis (H), and $x^{n}-a$, $x^{n}-a^{\prime}$ are irreducible over $\mathbb{Q}$. (For example, each primitive residue class modulo $p^{e+1}$ contains infinitely many squarefree integers.) Then $a\equiv a^{\prime}\pmod{p^{e}}$ but $a\not\equiv a^{\prime}\pmod{p^{e+1}}$. For odd $p$, the lifting-the-exponent lemma gives $v_{p}(a^{p-1}-1)=e$ and $v_{p}((a^{\prime})^{p-1}-1)\geq e+1$; for $p=2$ this is the same statement with $a^{p-1}-1=a-1$. Thus $r_{p}(a)=e-1$ while $r_{p}(a^{\prime})\geq e$, hence $d_{p}(a)=e-1$ and $d_{p}(a^{\prime})=e$. Let $m_{0}:=n-n/p^{e}$. By definition of $k_{p,m}$ we have $k_{p,m_{0}}=e-1$ but $k^{\prime}_{p,m_{0}}=e$, so $\mathbf{S}_{p}(a)\neq\mathbf{S}_{p}(a^{\prime})$. This proves (ii).
(iii) Global minimal period. Part (i) shows that for each $p\mid n$ the local shape depends only on $a\bmod p^{e_{p}+1}$. By the Chinese Remainder Theorem, the global shape depends only on $a\bmod\prod_{p^{e_{p}}\parallel n}p^{e_{p}+1}=M(n)$, so $M(n)$ is a period. Conversely, if $N$ is any modulus such that $a\equiv a^{\prime}\pmod{N}$ implies $\mathbf{S}(a)=\mathbf{S}(a^{\prime})$ for all $n$th–power–free $a,a^{\prime}$, then for each fixed $p\mid n$ it implies $\mathbf{S}_{p}(a)=\mathbf{S}_{p}(a^{\prime})$. By (ii) this forces $p^{e_{p}+1}\mid N$ for all $p\mid n$, hence $M(n)\mid N$. ∎
Remark 3.2 (Comparison with explicit periodicity).
In the explicit construction of integral bases for pure fields, Jakhar–Khanduja–Sangwan show that the pp–local exponents and corrections are functions of $d_{p}(a):=\min\{v_{p}(a^{p-1}-1)-1,e_{p}\}$ (for $v_{p}(a)=0$) and of $v_{p}(a)$, and that $d_{p}$ is stable under $a\bmod p^{e_{p}+1}$. This yields periodicity modulo $M(n)=\prod p^{e_{p}+1}$, and our Theorem 3.1 shows that this modulus is also minimal.
Example 3.3 (Nonic fields ($n=9$)).
Here $n=3^{2}$, so only $p=3$ contributes and the local period is $3^{e_{3}+1}=27$. For $3\nmid a$, put $r=v_{3}(a^{2}-1)-1$ and $d=\min\{r,2\}$. Then
$\begin{array}[]{c|c|c}a\!\!\!\pmod{27}&r&(k_{3,1},\dots,k_{3,8})\ \text{(nonzero entries)}\\ \hline\cr\pm 1\!\!\pmod{3}\ \not\equiv\pm 1\!\!\pmod{9}&0&\text{all }0\\ \pm 1\!\!\pmod{9}\ \not\equiv\pm 1\!\!\pmod{27}&1&k_{3,6}=k_{3,7}=k_{3,8}=1\\ \pm 1\!\!\pmod{27}&\geq 2&k_{3,6}=k_{3,7}=1,\ k_{3,8}=2\end{array}$
again matching the explicit basis in [JKS21].
Remark 3.4.
Separating the $p\mid n$ denominator profile from the factors $C_{m}(a)$ supported on primes dividing $a$ is standard in modern explicit constructions and is exactly the viewpoint in [JKS21]. For the restriction to primes $p\mid n$ in denominators see Remark 2.4, and for $n^{2}$–periodicity in small degrees see [GR17]. For the family $x^{p^{r}}-a$ and variants (with or without $(a,n)=1$), see [KCH24].
By Theorem 3.1, $\mathbf{S}_{p}(a)$ depends only on $a\bmod p^{e+1}$; there are only $p^{e+1}$ residue classes. For each class, the corresponding exponents $k_{p,m}$ and reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$ can be read off from the explicit construction of Jakhar–Khanduja–Sangwan [JKS21, Theorem 1.6 and Theorem 1.8]. It follows that there are only finitely many possible pp–local shapes, and the map
[…]
factors through $a\bmod p^{e+1}$.
As consequence, for each fixed $n$ there exists a finite table $\mathcal{T}(n)$ indexed by $a\bmod M(n)$ such that, for every $n$th–power–free $a$, the shape $\mathbf{S}(a)$ of an integral basis of $\mathcal{O}_{K_{a}}$ is read off from $\mathcal{T}(n)$ at the entry $a\bmod M(n)]$. The table is obtained by concatenating the local classes.
Theorem 3.5 (Computation table).
Fix $n$. There exists a finite lookup table $\mathcal{T}(n)$, indexed by the residue classes $a\bmod M(n)$, such that for any $n$th–power–free $a$ the shape of an integral basis of $\mathcal{O}_{K_{a}}$ is read off directly from the entry $a\bmod M(n)]\in\mathcal{T}(n)$. The table can be filled purely locally, prime–by–prime at $p\mid n$, using the explicit local recipes in [JKS21] for residue classes modulo $p^{e_{p}+1}$.
Proof.
By Theorem 3.1 the pp–part of the shape depends only on $a\bmod p^{e_{p}+1}$. Thus $\mathcal{T}(n)$ is the Cartesian product of the local tables at $p\mid n$ listing, for each residue class modulo $p^{e_{p}+1}$, the exponents $k_{p,m}$ and the reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$; these are exactly what the explicit construction in [JKS21, Theorem 1.6 and Theorem 1.8] provides. The global shape is obtained by Chinese remaindering of the local choices. ∎
We can develop relative version for the periodicity as follows.
Proposition 3.6 (Relative periodicity over tame bases).
Let $L/\mathbb{Q}$ be a number field with $\gcd(n,\Delta_{L})=1$, and let $K_{a,L}:=L(\sqrt[n]{a})$ with $a\in\mathbb{Z}$ $n$th–power–free. For each rational $p\mid n$ and each prime $\mathfrak{p}\mid p$ in $\mathcal{O}_{L}$, the $\mathfrak{p}$–local shape of an $\mathcal{O}_{L}$–integral basis of $K_{a,L}/L$ is determined by $a\bmod p^{e_{p}+1}$. Consequently the global relative shape is periodic modulo $M(n)$.
Proof.
Since $\gcd(n,\Delta_{L})=1$, every rational prime $p\mid n$ is unramified in $L$, so the local behavior at each prime $\mathfrak{p}\mid p$ is tame over $L$. The local denominator exponents and reduced correction terms in an $\mathcal{O}_{L}$–integral basis of $L(\sqrt[n]{a})/L$ at $\mathfrak{p}\mid p$ are governed by the same invariant $d_{p}(a)$ as in the absolute case (cf. [JKS21]), hence are determined by $a\bmod p^{e_{p}+1}$. Gluing over all $\mathfrak{p}\mid p$ and all $p\mid n$ yields periodicity modulo $M(n)$. ∎
The proof above mirrors the Kummer–cyclotomic situation where the decisive congruence lives modulo $(1-\zeta_{p})^{2}$, cf. [SMI21, Theorem. 1.1 and §6].
Example 3.7 (Relative periodicity over a quadratic tame base).
Let $n=6$ and $L=\mathbb{Q}(\sqrt{13})$. Then $\gcd(n,\Delta_{L})=1$ (here $\Delta_{L}=13$), so $2,3$ are unramified in $L$. In this field, $2$ is inert and $3$ splits:
$(2)=\mathfrak{p}_{2},\qquad(3)=\mathfrak{p}_{3,1}\,\mathfrak{p}_{3,2}.$
Consider $K_{a,L}:=L(\sqrt[6]{a})$ with $a\in\mathbb{Z}$ 66th–power–free. By Proposition 3.6, for each prime $\mathfrak{p}\mid 6$ the $\mathfrak{p}$–local shape of an $\mathcal{O}_{L}$–integral basis is determined by the class of $a$ modulo $p^{e_{p}+1}$ with $p\in\{2,3\}$ and $e_{2}=e_{3}=1$. Hence the global relative shape is periodic modulo
$M(6)=2^{\,2}\,3^{\,2}=36,$
and this modulus is minimal.
Explicit instance. Take $a\equiv 1\pmod{4}$ and $a\equiv 1\pmod{9}$ (e.g. $a=73$).
  • At $\mathfrak{p}_{2}\mid 2$ (inert): $d_{2}=1$, so
    $(k_{\mathfrak{p}_{2},1},k_{\mathfrak{p}_{2},2},k_{\mathfrak{p}_{2},3},k_{\mathfrak{p}_{2},4},k_{\mathfrak{p}_{2},5})=(0,0,1,1,1),$
    with reduced 22–adic corrections (mod $2$) as in the absolute sextic case:
    $\beta_{3,\mathfrak{p}_{2}}\equiv 1,\quad\beta_{4,\mathfrak{p}_{2}}\equiv\theta,\quad\beta_{5,\mathfrak{p}_{2}}\equiv\theta^{2}\pmod{2}$.
  • At each split prime $\mathfrak{p}_{3,i}\mid 3$ ($i=1,2$): $d_{3}=1$, so
    $(k_{\mathfrak{p}_{3,i},1},k_{\mathfrak{p}_{3,i},2},k_{\mathfrak{p}_{3,i},3},k_{\mathfrak{p}_{3,i},4},k_{\mathfrak{p}_{3,i},5})=(0,0,0,1,1),$
    with the same 33–adic reductions (mod $3$) for both $\mathfrak{p}_{3,1}$ and $\mathfrak{p}_{3,2}$:
    $\beta_{4,\mathfrak{p}_{3,i}}\equiv 1+\theta^{2},\qquad\beta_{5,\mathfrak{p}_{3,i}}\equiv\theta+\theta^{3}\pmod{3}$.
Thus, over $L$, the only effect of the base field is how many primes lie above each $p\mid 6$ (inert vs. split); the local exponents and reduced corrections at each $\mathfrak{p}$ are still decided entirely by
$a\bmod 4\quad\text{and}\quad a\bmod 9.$
In particular, the relative shape for $K_{a,L}/L$ depends only on $a\bmod 36$, and the four CRT cases from Example 2.8 carry over verbatim, with the 33–adic row duplicated on the two primes $\mathfrak{p}_{3,1},\mathfrak{p}_{3,2}$.
4. Applications of the minimal modulus
As in §2, when we write an integral basis in the $\{1,\theta,\dots,\theta^{n-1}\}$–coordinates we use the Jakhar–Khanduja–Sangwan form
{ $\Bigl\{\,1,\ \frac{\theta^{m}+\beta_{m}}{\;C_{m}(a)\,\prod_{p\mid n}p^{\,k_{p,m}}\;}\ :\ 1\leq m\leq n-1\Bigr\},$
where $C_{m}(a)$ is supported only at primes dividing $a$, while the pp–denominator exponents $k_{p,m}\geq 0$ and the linear corrections $\beta_{m}$ are determined locally at each $p\mid n$ from a finite amount of pp–adic information on $a$.
4.1. Sharp local index tests at primes $p\mid n$
We isolate two quick consequences for power–basis tests that fall directly out of the local classification.
Proposition 4.1 (Wieferich threshold when $e_{p}=1$).
Suppose $p\parallel n$ and $v_{p}(a)=0$ (equivalently, $(a,p)=1$). Then
$p\mid[\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]\quad\Longleftrightarrow\quad a^{p-1}\equiv 1\pmod{p^{2}}$.
Equivalently, $\mathbb{Z}[\theta]$ is pp–maximal if and only if $a^{p-1}\not\equiv 1\pmod{p^{2}}$.
Proof.
When $p\parallel n$ (so $e_{p}=1$) and $v_{p}(a)=0$, we have
$d_{p}(a)=\min\{\,v_{p}(a^{p-1}-1)-1,\,1\,\}\in\{0,1\}.$
If $a^{p-1}\not\equiv 1\pmod{p^{2}}$ then $v_{p}(a^{p-1}-1)=1$, so $d_{p}(a)=0$ and the explicit basis in Theorem 2.1 has no pp–power denominators; equivalently, $v_{p}([\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]])=0$. If $a^{p-1}\equiv 1\pmod{p^{2}}$ then $v_{p}(a^{p-1}-1)\geq 2$, so $d_{p}(a)=1$ and the explicit basis has a nontrivial pp–denominator, hence $v_{p}([\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]])>0$. ∎
In this paper we work in the fixed coordinates $\{1,\theta,\dots,\theta^{n-1}\}$ and in the normalized form where all factors supported on primes dividing $a$ are absorbed into the numerators (via the $C_{m}(a)$). In this normalization, the only primes that can occur in the denominators are those dividing $n$ (see Remark 2.4). Accordingly, Proposition 4.1 concerns the pp–part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$ at primes $p\mid n$; it is not meant to be a criterion involving primes dividing $a$ (indeed we assume $v_{p}(a)=0$ here).
Remark 4.2 (Index vs. monogeneity).
Proposition 4.1 is a criterion for the pp–adic contribution to the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, i.e. for pp–maximality of the power order $\mathbb{Z}[\theta]$. It does not decide whether $K_{a}$ is monogenic (whether there exists some $\alpha\in\mathcal{O}_{K_{a}}$ with $\mathcal{O}_{K_{a}}=\mathbb{Z}[\alpha]$).
Remark 4.3 (ABC and the Wieferich threshold at $p\parallel n$).
For $p\parallel n$ and $(a,p)=1$, our order–$1$ local shape at $p$ is decided by the Wieferich congruence $a^{p-1}\equiv 1\pmod{p^{2}}$ (Proposition 4.1). Assuming the abcabc conjecture, Silverman proved that for any fixed $A\geq 2$ there are $\gg\log x$ primes $p\leq x$ with $A^{p-1}\not\equiv 1\pmod{p^{2}}$; in particular there are infinitely many non–Wieferich primes for base $A$ [SIL88]. Consequently, for a fixed parameter $A$ and varying $n$, the order–$1$ local shape at $p\mid n$ is the pp–regular (no-denominator) branch for infinitely many $p$. Analogous non–Wieferich abundance over number fields under abcabc is known as well (see, e.g., Ding [DIN19]).
Proposition 4.4 (Local sharpness at every $p^{e}\parallel n$).
Fix $p^{e}\parallel n$ with $e\geq 1$ and assume $(a,p)=1$. There are integers
$a\equiv 1+p^{e}u\not\equiv 1\pmod{p^{e+1}}\quad\text{and}\quad a^{\prime}\equiv 1+p^{e+1}u^{\prime}\pmod{p^{e+2}}$
($u,u^{\prime}\in\mathbb{Z}$, $p\nmid uu^{\prime}$) such that the pp–parts of the shapes of $K_{a}$ and $K_{a^{\prime}}$ differ. Equivalently, the local modulus $p^{e+1}$ is minimal.
Proof.
For odd $p$, the LTE lemma gives $v_{p}\bigl((1+p^{e}u)^{p-1}-1\bigr)=e$, hence $r_{p}(a)=e-1$, while $r_{p}(a^{\prime})\geq e$. Thus $d_{p}(a)=e-1<e=d_{p}(a^{\prime})$ and the local denominator pattern changes with $a\bmod p^{e+1}$. For $p=2$ the same conclusion holds with $r_{2}(a)=v_{2}(a-1)-1$ by direct expansion. The resulting change of $k_{p,m}$ (and of the reduced corrections $\beta_{m}\bmod p^{k_{p,m}}$) is built into the Jakhar–Khanduja–Sangwan explicit bases. ∎
Write $p^{e}\parallel n$ and $n_{p}:=n/p^{e}$. The Jakhar–Khanduja–Sangwan discriminant formula shows that $v_{p}(d_{K_{a}})$ depends only on $e$ and $t_{p}:=\min\{r_{p}(a),e\}$. Specialising that formula, we obtain the following exact step size.
Proposition 4.5 (Exact jump in $v_{p}(d_{K_{a}})$ across one local step).
Suppose $t,t^{\prime}\in\{0,1,\dots,e\}$ with $t^{\prime}=t+1$. There exist $n$th–power–free integers $a,a^{\prime}$ with $a^{\prime}\equiv a\pmod{p^{e}}$, $t_{p}(a)=t$ and $t_{p}(a^{\prime})=t^{\prime}$ such that
$v_{p}\!\bigl(d_{K_{a^{\prime}}}\bigr)\;=\;v_{p}\!\bigl(d_{K_{a}}\bigr)\;-\;2\,n_{p}\,p^{\,e-(t+1)}.$
In particular, $v_{p}(d_{K_{a}})$ is constant on $a\bmod p^{e+1}$–classes and changes by the displayed discrete steps across $a\bmod p^{e}$–classes.
Proof.
The discriminant exponent has the form
$v_{p}(d_{K_{a}})=\begin{cases}n\,e-2\,n_{p}\displaystyle\sum_{j=1}^{t}p^{\,e-j},&t\geq 1,\\[4.0pt] n\,e,&t\leq 0,\end{cases}$
so raising $t$ to $t+1$ subtracts precisely $2\,n_{p}\,p^{\,e-(t+1)}$. The existence of such $a,a^{\prime}$ with fixed residue modulo $p^{e}$ and $t_{p}$ differing by one is ensured by Proposition $4.$ (We are specializing [JKS21, Theorem. 1.A] to a fixed $p^{e}\,\|\,n$.)
4.2. Counting and density of shape classes
By Theorem 3.1, for fixed $n$ the (global) shape $\mathbf{S}(a)$ depends only on the residue class of $a$ modulo $M(n)$. Therefore, any question of the form “how often does a given shape (or a union of shapes) occur as $|a|\leq X$?” reduces to counting those integers $a$ with $|a|\leq X$ that lie in prescribed congruence classes modulo $M(n)$ and are $n$th-power-free (because our fields $K_{a}=\mathbb{Q}(\sqrt[n]{a})$ are only defined for $n$th-power-free parameters). This is the analytic input behind our later shape density statements. Recall that an integer $a\neq 0$ is $n$th-power-free if no prime $p$ satisfies $p^{n}\mid a$. Equivalently, in the prime factorization $|a|=\prod_{p}p^{\alpha_{p}}$ one has $\alpha_{p}<n$ for every $p$. The sign of $a$ plays no role for $n$th-power-freeness.
We use a standard counting function for $n$th–power–free integers in arithmetic progressions. Fix integers $q\geq 1$ and r(modq)r\ ($\mathrm${mod}\ q). We define
$\mathcal{F}_{n}(X;q,r):=\#\bigl\{\,a\in\mathbb{Z}:\ 0<|a|\leq X,\ a\equiv r\ (\mathrm{mod}\ q),\ a\ \text{is $n$th-power-free}\,\bigr\}.$
Thus $\mathcal{F}_{n}(X;q,r)$ counts $n$th–power–free integers in the arithmetic progression $a\equiv r\pmod{q}$, up to height $X$ (counting both positive and negative integers; this explains the factor $2X$ that will appear in main terms). If $\mathcal{F}_{n}(X;q,r)$ grows linearly with $X$, then the leading constant is exactly the natural density (with respect to height $|a|$) of $n$th–power–free integers in that progression. Concretely, whenever the limit exists,
$\delta_{n}(q,r)\ :=\ \lim_{X\to\infty}\frac{\mathcal{F}_{n}(X;q,r)}{2X}$
is the proportion of integers (ordered by $|a|$) that lie in the class $r\bmod q$ and are $n$th–power–free. Theorem 4.8 computes this density and makes explicit how it depends on $r$.
Not every residue class modulo $q$ can contain infinitely many $n$th–power–free integers. For example, if $q$ contains a prime power $p^{\alpha}$ with $\alpha\geq n$ and $r\equiv 0\pmod{p^{\alpha}}$, then every $a\equiv r\pmod{q}$ is divisible by $p^{n}$, hence cannot be $n$th–power–free (except possibly $a=0$, which is excluded by $n$th–power–freeness). The following definitions isolate precisely this obstruction
Definition 4.6 (Admissibility notions).
Fix $q\geq 1$ and $n\geq 2$. For a residue $r\bmod q$ and a prime power $p^{\alpha}\parallel q$ set $v_{p}(r)$ to be the pp–adic valuation of a chosen representative of $r$.
  • (a)
    We say that $r$ is nn–admissible (weak form) if for every $p^{\alpha}\parallel q$ with $\alpha\geq n$ we have $v_{p}(r)<n$. Equivalently, the progression $a\equiv r\pmod{q}$ does not force the divisibility $p^{n}\mid a$ for any prime $p$.
  • (b)
    We say that $r$ is strictly nn–admissible if for every $p^{\alpha}\parallel q$ we have $v_{p}(r)<\min\{\alpha,n\}$. This is stronger than (a) and will be exactly the condition that makes the main term independent of $r$.
Remark 4.7 (Terminology).
Definition 4.6 concerns $n$th–power–free integers in a general arithmetic progression modulo $q$. Later, for the specific modulus $q=M(n)$ coming from the local theory, we will often restrict to the smaller set of residue classes with $v_{p}(r)\leq 1$ for every $p\mid n$; to avoid confusion we will call these shape–admissible classes (Definition 4.11 below).
For a nonzero integer $a$, the indicator of $n$th–power–freeness is
$1_{\text{$n$--free}}(a)\ :=\begin{cases}1,&\text{if $p^{n}\nmid a$ for every prime $p$,}\\ 0,&\text{otherwise.}\end{cases}$
A classical identity (see, e.g., [TEN15, Ch. III]) is
(2) $1_{\text{$n$--free}}(a)\ =\ \sum_{d^{n}\mid a}\mu(d).$
Here $\mu$ is the Möbius function. One quick justification is to factor the right-hand side as an Euler product: since $\mu(d)=0$ unless $d$ is squarefree, we have
$\sum_{d^{n}\mid a}\mu(d)=\prod_{p}\Bigl(1+\mu(p)\,\mathbf{1}_{\{p^{n}\mid a\}}\Bigr)=\prod_{p}\Bigl(1-\mathbf{1}_{\{p^{n}\mid a\}}\Bigr),$
which equals $1$ if no prime power $p^{n}$ divides $a$, and equals 0 otherwise, exactly as desired. Our counting function $\mathcal{F}_{n}(X;q,r)$ ignores $a=0$ anyway, since 0 is not $n$th–power–free.
We now evaluate $\mathcal{F}_{n}(X;q,r)$ asymptotically, keeping track of the precise dependence on $r$.
Theorem 4.8 (Equidistribution with explicit $r$-dependence).
Fix $q\geq 1$, $n\geq 2$, and a residue class r(modq)r\ ($\mathrm${mod}\ q). Then, as $X\to\infty$,
(3) $\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\cdot\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}\,\prod_{p^{\alpha}\parallel q}\Bigl(1-\mathbf{1}_{\{v_{p}(r)\geq\min(\alpha,n)\}}\,p^{\min(\alpha,n)-n}\Bigr)\;+\;O\!\bigl(X^{1/n}\bigr).$
In particular:
  • If $r$ is not $n$-admissible, then $\mathcal{F}_{n}(X;q,r)=O(1)$.
  • If $r$ is strictly $n$-admissible, then the $r$-dependent product in (3) equals $1$, and the main term simplifies to
    $\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\cdot\frac{1}{\zeta(n)}\,\prod_{p\mid q}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}\;+\;O\!\bigl(X^{1/n}\bigr),$
    independently of $r$.
Proof.
By definition,
$\mathcal{F}_{n}(X;q,r)=\sum_{\begin{subarray}{c}0<|a|\leq X\\ a\equiv r\ (\mathrm{mod}\ q)\end{subarray}}1_{\text{$n$-free}}(a).$
It follows that
$\mathcal{F}_{n}(X;q,r)=\sum_{\begin{subarray}{c}0<|a|\leq X\\ a\equiv r\ (\mathrm{mod}\ q)\end{subarray}}\ \sum_{d^{n}\mid a}\mu(d).$
The inner condition $d^{n}\mid a$ forces $d\leq X^{1/n}$ (since $0<|a|\leq X$), so the double sum is finite; there is no convergence issue, and we may swap the order of summation:
$\mathcal{F}_{n}(X;q,r)=\sum_{1\leq d\leq X^{1/n}}\mu(d)\cdot\#\bigl\{\,0<|a|\leq X:\ a\equiv r\ (\mathrm{mod}\ q),\ d^{n}\mid a\,\bigr\}.$
Fix $d\leq X^{1/n}$. We must count $a$ with
$a\equiv r\ (\mathrm{mod}\ q)\qquad\text{and}\qquad a\equiv 0\ (\mathrm{mod}\ d^{n}).$
Let
$g:=\gcd(q,d^{n}),\qquad L:=\mathrm{lcm}(q,d^{n})=\frac{q\,d^{n}}{g}.$
A standard compatibility criterion says that the two congruences are simultaneously solvable if and only if $r\equiv 0\ (\mathrm{mod}\ g)$, i.e. $g\mid r$. Indeed, any solution must be congruent to 0 modulo $g$, hence $r\equiv a\equiv 0\ (\mathrm{mod}\ g)$. Conversely, if $g\mid r$, then dividing the second congruence by $g$ reduces it to a single class modulo $q/g$, and since $\gcd(q/g,\,d^{n}/g)=1$, the Chinese Remainder Theorem produces a unique solution modulo $L$. Hence there is either:
  • no solution if $g\nmid r$, or
  • exactly one residue class $a\equiv a_{0}\ (\mathrm{mod}\ L)$ if $g\mid r$.
In the solvable case, the number of integers $a$ with $0<|a|\leq X$ in one fixed class modulo $L$ is
$\frac{2X}{L}+O(1),$
since such integers form a two-sided arithmetic progression with step $L$. Therefore
#{ $\bigl\{\,0<|a|\leq X:\ a\equiv r\ (\mathrm{mod}\ q),\ d^{n}\mid a\,\bigr\}=\mathbf{1}_{\{g\mid r\}}\Bigl(\frac{2X}{L}+O(1)\Bigr)=\mathbf{1}_{\{g\mid r\}}\Bigl(\frac{2X}{q}\cdot\frac{g}{d^{n}}+O(1)\Bigr).$
Plugging this into the $d$-sum yields
$\mathcal{F}_{n}(X;q,r)=\frac{2X}{q}\sum_{1\leq d\leq X^{1/n}}\mu(d)\,\frac{g}{d^{n}}\,\mathbf{1}_{\{g\mid r\}}\;+\;O\!\Bigl(\sum_{d\leq X^{1/n}}1\Bigr).$
Since $\sum_{d\leq X^{1/n}}1\ll X^{1/n}$, the error term is $O(X^{1/n})$.
Now we analyze
$S(q,r):=\sum_{d\geq 1}\mu(d)\,\frac{\gcd(d^{n},q)}{d^{n}}\,\mathbf{1}_{\{\gcd(d^{n},q)\mid r\}}.$
Only squarefree $d$ contribute because $\mu(d)=0$ otherwise, so write $d=\prod_{p}p^{\nu_{p}(d)}$ with $\nu_{p}(d)\in\{0,1\}$. Also write $q=\prod_{p}p^{\alpha_{p}}$ where $\alpha_{p}=v_{p}(q)$. Then
$\gcd(d^{n},q)=\prod_{p}p^{\min(n\nu_{p}(d),\alpha_{p})}.$
Consequently, the condition $\gcd(d^{n},q)\mid r$ is equivalent to requiring, for every prime $p$ with $\nu_{p}(d)=1$, that
$v_{p}(r)\geq\min(n,\alpha_{p}).$
Moreover,
$\frac{\gcd(d^{n},q)}{d^{n}}=\prod_{p:\,\nu_{p}(d)=1}p^{\min(n,\alpha_{p})-n}.$
This shows that the sum $S(q,r)$ factors as a product of independent local contributions over primes. For a given prime $p$, only $\nu=0$ and $\nu=1$ occur:
$\sum_{\nu\in\{0,1\}}\mu(p^{\nu})\,p^{\min(n\nu,\alpha_{p})-n\nu}\,\mathbf{1}_{\{\min(n\nu,\alpha_{p})\leq v_{p}(r)\}}=1-\mathbf{1}_{\{v_{p}(r)\geq\min(n,\alpha_{p})\}}\,p^{\min(n,\alpha_{p})-n}.$
If $p\nmid q$ then $\alpha_{p}=0$ and the local factor becomes $1-p^{-n}$.
Therefore,
$S(q,r)=\prod_{p\nmid q}(1-p^{-n})\cdot\prod_{p^{\alpha}\parallel q}\Bigl(1-\mathbf{1}_{\{v_{p}(r)\geq\min(\alpha,n)\}}\,p^{\min(\alpha,n)-n}\Bigr).$
Since $\prod_{p}(1-p^{-n})=\zeta(n)^{-1}$, we may rewrite
$\prod_{p\nmid q}(1-p^{-n})=\frac{1}{\zeta(n)}\prod_{p\mid q}(1-p^{-n})^{-1}.$
Substituting back gives precisely the main term in (3).
If $r$ is not $n$-admissible, then there exists $p^{\alpha}\parallel q$ with $\alpha\geq n$ and $v_{p}(r)\geq n$, so every $a\equiv r\pmod{q}$ satisfies $p^{n}\mid a$, hence there are no $n$th-power-free integers in the progression for $|a|$ large (so $\mathcal{F}_{n}(X;q,r)=O(1)$).
If $r$ is strictly $n$-admissible, then $v_{p}(r)<\min(\alpha,n)$ for every $p^{\alpha}\parallel q$, so all indicators in the $r$-dependent product vanish, and the main term is independent of $r$. ∎
Dividing (3) by $2X$ and letting $X\to\infty$ shows that the density $\delta_{n}(q,r)$ exists for every residue class $r$, and is given by the constant multiplying $\frac{2X}{q}$. This is exactly what we need later: once a shape (or a union of shapes) is specified by congruence conditions on $a\pmod{M(n)}$, its natural density among $n$th-power-free parameters follows by summing these constants over the relevant residue classes.
Remark 4.9.
Assume $p\parallel n$ and $(a,p)=1$. Proposition 4.1 gives a criterion for the $p$-part of the index $\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, governed by the congruence $a^{p-1}\equiv 1\pmod{p^{2}}$. Among units modulo $p^{2}$, this congruence holds for a proportion $1/p$. Therefore a positive proportion of parameters $a$ (subject to the $n$th-power-free constraint) satisfy $p\mid[\mathcal{O}_{K_{a}}:\mathbb{Z}[\theta]]$, i.e. $\theta$ does not generate $\mathcal{O}_{K_{a}}$. This does not decide whether $K_{a}$ is monogenic via some other generator.
Corollary 4.10.
If $q$ is nn–full, i.e. $v_{p}(q)\geq n$ for every $p\mid q$, then all nn–admissible residues r(modq)r\ ($\mathrm${mod}\ q) are strictly nn–admissible, and therefore the $n$th–power–free integers are equidistributed among the nn–admissible classes with common density
$\frac{1}{q}\,\zeta(n)^{-1}\,\prod_{p\mid q}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}.$
In our applications we take $q=M(n)=\prod_{p^{e}\parallel n}p^{e+1}$. Since typically $e+1<n$, Corollary 4.10 need not apply, and the uniform density across all admissible classes mod $M(n)$ may fail. Theorem 4.8, however, still yields the correct main term with the explicit rr–dependence. For many counting questions it suffices to restrict attention to strictly nn–admissible classes (e.g. those with $v_{p}(r)\in\{0,1\}$ at $p\mid n$), in which case uniformity is restored.
Definition 4.11 (Shape–admissible residue classes modulo $M(n)$).
A residue class r(modM(n))r\ ($\mathrm${mod}\ M(n)) is called shape–admissible if
$v_{p}(r)\leq 1\qquad\text{for every prime }p\mid n.$
Note that all denominator phenomena are localised at $p\mid n$, and the shape is constant on each residue class modulo $M(n)$. Hence the shape map
$\mathsf{Sh}_{n}:\ \{a\in\mathbb{Z}\text{ $n$th–power–free}\}\ \longrightarrow\ \mathcal{S}_{n}$
factors through $a\bmod M(n)$. Fix any union $\mathcal{R}\subseteq(\mathbb{Z}/M(n)\mathbb{Z})$ of shape–admissible classes (in the sense of Definition 4.11). Let
$\mathcal{N}_{\mathcal{R}}(X)\ :=\ \#\{\,|a|\leq X:\ a\text{ $n$th–power–free and }a\bmod M(n)\in\mathcal{R}\,\}.$
Theorem 4.12 (Equidistribution and density).
As $X\to\infty$,
$\mathcal{N}_{\mathcal{R}}(X)\ =\ \frac{2X}{M(n)}\,\biggl(\prod_{p}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)\biggr)\,\prod_{p\mid n}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}\,\#\mathcal{R}\ +\ o(X).$
In particular, the natural density of $n$th–power–free $a$ with $a\bmod M(n)\in\mathcal{R}$ exists and equals
$\delta(\mathcal{R})\ =\ \frac{\#\mathcal{R}}{M(n)}\cdot\frac{1}{\zeta(n)}\prod_{p\mid n}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}.$
Proof.
Write the indicator of $n$th–power–freeness as $1_{\text{$n$–free}}(a)=\sum_{d^{n}\mid a}\mu(d)$ and count by inclusion–exclusion in the progression $a\equiv r\pmod{M(n)}$. If $r$ is shape–admissible, then $(d,M(n))=1$ in all contributing terms; the main term is
$\frac{2X}{M(n)}\sum_{(d,M(n))=1}\frac{\mu(d)}{d^{n}}=\frac{2X}{M(n)}\,\prod_{p\nmid M(n)}\Bigl(1-\frac{1}{p^{n}}\Bigr)=\frac{2X}{M(n)}\,\prod_{p}\Bigl(1-\frac{1}{p^{n}}\Bigr)\prod_{p\mid n}\Bigl(1-\frac{1}{p^{\,n}}\Bigr)^{-1}.$
Summing over $r\in\mathcal{R}$ gives the formula; standard partial summation bounds yield $o(X)$.Summing over $r\in\mathcal{R}$ gives the formula; standard partial summation bounds yield $o(X)$.
This is the standard Möbius‑inversion count for $k$-free numbers in arithmetic progressions (see [TEN15]), specialized to $k=n$ and the modulus $M(n)$ singled out by the local theory; the main term factors as a product of local proportions, matching the CRT‑gluing of local shapes. For quantitative refinements (error terms for $k$-free in APs), see Nunes [NUN22] and Parry [PAR21].
Remark 4.13 (Local factorisation of the count).
By the Chinese Remainder Theorem, $\mathcal{R}$ factors as $\mathcal{R}=\prod_{p\mid n}\#\mathcal{R}_{p}$, where $\mathcal{R}_{p}$ is the chosen set of local residues modulo $p^{e_{p}+1}$ realising the desired pp–component of the shape. Consequently the global density $\delta(\mathcal{R})$ factorises as a product of local densities.
Example 4.14 (Counting and density for $n=4$).
Here $M(4)=8$ and the shape–admissible residue classes (in the sense after Theorem 4.12) are those r(mod 8)r\ ($\mathrm${mod}\ 8) with $v_{2}(r)\leq 1$, i.e. $r\in\{1,3,5,7,2,6\}$ (six classes). The three quartic shapes are:
$\begin{array}[]{c|c}\text{shape }(k_{2,1},k_{2,2},k_{2,3})&\text{classes }r\ (\mathrm{mod}\ 8)\\ \hline\cr(0,1,2)&\{1\}\\ (0,1,1)&\{5\}\\ (0,0,0)&\{3,7,2,6\}\end{array}$
By Theorem 4.12, for any union $\mathcal{R}$ of shape–admissible classes one has
$\delta(\mathcal{R})=\frac{\#\mathcal{R}}{8}\cdot\frac{1}{\zeta(4)}\Bigl(1-\frac{1}{2^{4}}\Bigr)^{-1}.$
Hence the individual shape densities are
$\delta_{(0,1,2)}=\delta_{(0,1,1)}=\frac{1}{8}\cdot\frac{1}{\zeta(4)}\Bigl(1-\frac{1}{2^{4}}\Bigr)^{-1}=\frac{12}{\pi^{4}},\qquad\delta_{(0,0,0)}=\frac{4}{8}\cdot\frac{1}{\zeta(4)}\Bigl(1-\frac{1}{2^{4}}\Bigr)^{-1}=\frac{48}{\pi^{4}}.$
Example 4.15 (Counting and density for $n=6$).
[…]
Write
$C_{6}\ :=\ \frac{1}{\zeta(6)}\Bigl(1-\frac{1}{2^{6}}\Bigr)^{-1}\Bigl(1-\frac{1}{3^{6}}\Bigr)^{-1}.$
Using the CRT rows in Example 2.8, the four sextic shape‑types occur on the following numbers of classes:
[…]
Therefore, by Theorem 4.12,
$\delta_{\text{both}}=\frac{2}{36}\,C_{6}=\frac{1}{18}\,C_{6},\quad\delta_{\text{only }2}=\frac{6}{36}\,C_{6}=\frac{1}{6}\,C_{6},\quad\delta_{\text{only }3}=\frac{4}{36}\,C_{6}=\frac{1}{9}\,C_{6},\quad\delta_{\text{none}}=\frac{12}{36}\,C_{6}=\frac{1}{3}\,C_{6}.$
(These four densities sum to $\tfrac{24}{36}C_{6}$, reflecting that we have restricted to the 2424 shape–admissible classes.)
When $(a,p)=1$, the pp–local behaviour is governed by
$r_{p}(a):=v_{p}(a^{p-1}-1)-1,\qquad d_{p}(a):=\min\{r_{p}(a),e_{p}\}.$
The following easy count (for odd $p$) describes $r_{p}$ among units modulo $p^{e+1}$.
Lemma 4.16 (Distribution of $r_{p}$ among units).
Let $p$ be odd and $e\geq 1$. Among $u\in(\mathbb{Z}/p^{e+1}\mathbb{Z})^{\times}$ one has
$\mathbb{P}\!\left(v_{p}(u^{p-1}-1)\geq k\right)=p^{-(k-1)}\quad(k\geq 1),$
and hence, for $0\leq j<e$,
$\mathbb{P}(r_{p}=j)=\frac{p-1}{p^{\,j+1}},\qquad\mathbb{P}(r_{p}\geq e)=\frac{1}{p^{\,e}}.$
Proof.
$(\mathbb{Z}/p^{t}\mathbb{Z})^{\times}$ is cyclic of order $p^{t-1}(p-1)$. The congruence $x^{p-1}\equiv 1\pmod{p^{t}}$ has exactly p1p-1 solutions. Fix $k\geq 1$. The condition $v_{p}(u^{p-1}-1)\geq k$ is equivalent to $u^{p-1}\equiv 1\pmod{p^{k}}$, so $u\bmod p^{k}$ must lie in one of these p1p-1 solution classes. Each such class modulo $p^{k}$ has $p^{e+1-k}$ lifts to a class modulo $p^{e+1}$, hence the number of units $u\ (\bmod\ p^{e+1})$ with $v_{p}(u^{p-1}-1)\geq k$ is $(p-1)p^{e+1-k}=(p-1)p^{e-(k-1)}$, giving the stated probabilities; the formulae follow. The case $p=2$ is similar with the usual structure of $(\mathbb{Z}/2^{t}\mathbb{Z})^{\times}$. ∎
Corollary 4.17 (Wieferich split when $e_{p}=1$).
If $p\parallel n$ and $(a,p)=1$, then among the pp–admissible classes modulo $p^{2}$ the proportion with $r_{p}=0$ (i.e. $v_{p}(a^{p-1}-1)=1$) is $1-\frac{1}{p}$, and the proportion with $r_{p}\geq 1$ is $\frac{1}{p}$.
Proof.
Apply Lemma 4.16 with $e=1$. ∎